# FRISM: Fine-Grained Reasoning Injection via Subspace-Level Model Merging for Vision–Language Models

### v1

@@ Abstract
Efficiently enhancing the reasoning capabilities of Vision-Language Models (VLMs) by merging them with Large Reasoning Models (LRMs) has emerged as a promising direction. However, existing methods typically operate at a coarse-grained layer level, which often leads to a trade-off between injecting reasoning capabilities and preserving visual capabilities. To address this limitation, we propose FRISM (Fine-grained Reasoning Injection via Subspace-level model Merging), a fine-grained reasoning injection framework based on subspace-level model merging. Observing that reasoning capabilities are encoded in distinct subspaces, FRISM decomposes LRM task vectors via Singular Value Decomposition (SVD) and adaptively tunes the scaling coefficients of each subspace through learning to realize fine-grained reasoning injection. Furthermore, we introduce a label-free self-distillation learning strategy with a dual-objective optimization using common vision-language perception datasets. Extensive experiments demonstrate that FRISM effectively improves reasoning capabilities without compromising the model’s original visual capabilities by consistently achieving state-of-the-art performance across diverse visual reasoning benchmarks.
Machine Learning, ICML
cyhuang24@m.fudan.edu.cn
@@ 1 Introduction
In recent years, represented by LLaVA (Liu et al., 2023, 2024), Qwen-VL (Wang et al., 2024c; Bai et al., 2025), and InternVL (Chen et al., 2024b; Zhu et al., 2025), Vision Language Models (VLMs) built upon Large Language Models (LLMs) have achieved remarkable success, demonstrating strong performance across a wide range of vision language tasks, including image recognition, visual understanding, and text generation (Liu et al., 2023, 2024; Cao et al., 2024; Brooks et al., 2023). In parallel, the emergence of large language reasoning models (LRMs) like Deepseek-R1 (DeepSeek-AI, 2025) and OpenAI-o1 (Jaech et al., 2024), has significantly advanced the frontier of model capabilities, especially in complex tasks like mathematical reasoning, problem solving, programming, and verification (Chen et al., 2025a; Zheng et al., 2025). Accordingly, many recent studies have focused on improving the reasoning capabilities of VLMs (Wang et al., 2025c), proposing various benchmarks (Yang et al., 2025; Wang et al., 2024b) and training strategies (Wang et al., 2025a; Huang et al., 2025). These works consistently show the effectiveness of enhancing reasoning capabilities to improving the performance of VLMs.
@@
Due to the lack of labeled vision-language reasoning training datasets and the heavy post-training costs, efficiently equipping VLMs with reasoning capabilities has become an important research direction. Among existing methods (Xiao et al., 2025; Huang et al., 2025), model merging achieves this by integrating the weights of LRMs and VLMs, incurring minimal training overhead and data annotation costs. Bring Reason to Vision (BR2V) (Chen et al., 2025b) first demonstrates the feasibility of reasoning injection from LRMs to VLMs through model merging. However, common model merging techniques are developed for multi-task settings, which may compromise the visual capabilities of the merged models and cause them to fail. To this end, FRANK (Wei and Chen, 2025) proposes a layer-wise model merging method by identifying shallow and deeper layers and employing a Taylor-derived closed-form solution to guide the merging process. IP-Merging (Hu et al., 2025) only merges a subset of typical layers based on the similarity between the weights of VLMs and LRMs while leaving the remaining layers unchanged. Although effective to some extent, these methods still exhibit a significant performance gap compared to post-training-based reasoning VLMs.
@@
$W_{\text{M}}=W_{\text{base}}+\sum_{m\in\left[\text{vlm},\text{lrm}\right]}\lambda_{m}\cdot\tau_{m},$ (1)
where $W_{\text{M}}$ represents the weights of the merged layer, and $W_{\text{base}}$ denotes the weights of the pre-trained base model. The term $\tau_{m}$ refers to the task vector derived from model $m$, and $\lambda_{m}$ is the associated merging coefficient, where $m\in\{\text{vlm},\text{lrm}\}$ corresponds to VLM and LRM, respectively. Fig. 1 further shows the diagram of (a) individual models, (b) layer-wise merging. For individual models, $\lambda_{\text{vlm}}=1,\lambda_{\text{lrm}}=0$ recovers the original VLM, while $\lambda_{\text{vlm}}=0,\lambda_{\text{lrm}}=1$ recovers the original LRM. Layer-wise merging methods assign layer-dependent values to $\lambda_{\text{vlm}}$ and $\lambda_{\text{lrm}}$, where a larger $\lambda_{\text{vlm}}$ indicates stronger visual capability and a larger $\lambda_{\text{lrm}}$ indicates stronger reasoning capability. However, such layer-wise balancing is inherently coarse-grained and can only produce a compromise: tuning $\lambda_{\text{vlm}}$ and $\lambda_{\text{lrm}}$ at the layer level entangles the two capabilities, leading to unavoidable losses in either vision or reasoning capabilities. Fig. 2 shows the average performance on two vision-language (VL) reasoning datasets, MMStar (Chen et al., 2024a) and R1-OneVision (Yang et al., 2025), and two VL perception datasets, TextVQA (Singh et al., 2019) and POPE (Li et al., 2023b). The results show that existing methods usually lead to a trade-off between VL reasoning and perception performance.
In this paper, we observe that a layer is not the atomic unit of model capability: even within a single layer, reasoning capabilities and visual representations can reside in distinct subspaces. Motivated by the low-rank nature of parameter updates in LLMs (Cai et al., 2025; Ping et al., 2024), we use the subspaces of the LRM task vectors obtained through singular value decomposition (SVD) as the reasoning prior and tune the injection magnitude of each subspace. We find that the gains from different ranks peak at different magnitudes. This heterogeneity suggests that a uniform layer-wise scaling coefficient is insufficient: it inevitably couples beneficial reasoning, which may require high magnitude, with visual noise, thereby limiting overall improvements.
Thus, to achieve the maximum reasoning capabilities with the minimum degree of vision loss, we propose FRISM (Fine-grained Reasoning Injection via Subspace-level model Merging). We change the merging paradigm to:
@@
  • We rethink the paradigm of merging VLMs and LRMs to efficiently enhance visual reasoning capabilities. Instead of coarse-grained merging in the layer-level, we propose a fine-grained merging paradigm to realize more effective reasoning injection.
  • We first propose a subspace-level model merging for reasoning injection from LRMs to VLMs. By decomposing LRM task vectors via SVD and modulating at the subspace granularity, our method enables the adaptive amplification of reasoning-dominant components while preserving the visual manifold.
  • We design label-free self-distillation training to relieve the scarcity of multimodal reasoning data. Through a dual-objective optimization that safeguards visual consistency while maximizing spectral injection intensity, our approach achieves stable and efficient reasoning transfer without requiring VL reasoning dataset.
@@ 2 Related Work
Model Merging integrates models finetuned on different tasks within the parameter space to assimilate diverse capabilities (Yang et al., 2024b). In multi-task settings, the critical challenge for model merging is weight interference. Various methods are proposed to address this issue, including calculating information matrices (Jin et al., 2022; Matena and Raffel, 2022), merging coefficient tuning (Ilharco et al., 2022; Yang et al., 2023), task vector pruning (Yadav et al., 2023; Yu et al., 2024; Huang et al., 2024; Gargiulo et al., 2025; Marczak et al., 2025), and dynamic routing (Ye et al., 2025; Tang et al., 2024; Lu et al., 2024b; Shen et al., 2024; Yang et al., 2024c), Recently, this paradigm has been extended to merge VLMs and LRMs (Chen et al., 2025b), aiming to transfer the reasoning capabilities between models trained on different modalities. However, methods designed under multi-task settings often incur severe performance loss. To relieve this issue, FRANK (Wei and Chen, 2025) performs layer-wise weighted fusion based on Taylor expansion and IP-Merging (Hu et al., 2025) mitigates cross-modal conflicts via parameter projection based on weight similarity. Nevertheless, limited by coarse granularity or static thresholding, these methods struggle to distinguish parameters beneficial for reasoning from those that cause visual interference. Therefore, effective Reasoning-to-Vision transfer requires not merely conflict mitigation but also precise control over task vectors. To this end, we propose FRISM, which fine-grainedly localizes and preserves parameter components that are critical for reasoning without harming visual capabilities.
Subspace Decomposition technique aims to factorize high-dimensional matrices to reveal latent structures. Methods such as Non-negative Matrix Factorization (NMF) and Independent Component Analysis (ICA) demonstrate advantages in interpretability or disentanglement. Among them, SVD serves as a fundamental subspace decomposition technique offering optimal low-rank approximations. It stands as a cornerstone in deep learning for model compression (Idelbayev and Carreira-Perpinán, 2020) and parameter-efficient fine-tuning (PEFT) (Hu et al., 2022; Meng et al., 2024). Recent studies (Cai et al., 2025; Sharma et al., 2024; Zhang et al., 2023; Ping et al., 2024) indicate that model capabilities are not uniformly distributed but are distinctively encoded within specific rank subspaces. Motivated by these findings, we treat LRM task vectors as inference priors. By recalibrating their spectral distribution via SVD, FRISM seamlessly adapts textual reasoning priors to the VL domain, thereby enhancing VLM reasoning capabilities without compromising visual performance.
@@ 3 Method
where $\tau$ are task vectors. Existing methods usually conduct model-level merging (Chen et al., 2025b):
$\theta_{\text{vlr}}=\theta_{\text{base}}+\lambda_{\text{vlm}}\cdot\tau_{\text{vlm}}+\lambda_{\text{lrm}}\cdot\tau_{\text{lrm}},$ (4)
@@
$\theta_{\text{vlr}}^{(l)}=\theta_{\text{base}}^{(l)}+\lambda_{\text{vlm}}^{(l)}\cdot\tau_{\text{vlm}}^{(l)}+\lambda_{\text{lrm}}^{(l)}\cdot\tau_{\text{lrm}}^{(l)},\quad l=1..L$ (5)
@@
Motivation. We hypothesize that reasoning capabilities and visual representations reside in distinct subspaces within the parameter space. Consequently, a coarse-grained, layer-wise merging coefficient may be insufficient to disentangle these capabilities. To validate this hypothesis, we analyze the impact of merging different subspaces of the LRM task vector. Specifically, we decompose the task vector of DeepSeek-R1-Distill-Qwen-7B, via SVD and inject specific rank subspaces into the VLM, Qwen2.5-VL-Instruct, with varying scaling coefficients. We observe that (1) different subspaces achieve peak performance at distinctly different scaling coefficients; (2) standard layer-wise merging results in suboptimal performance. These observations suggest that reasoning capabilities are encoded in specific subspaces. Therefore, to achieve maximum reasoning capabilities with minimal vision loss, it is essential to shift from coarse-grained layer-level merging to a fine-grained subspace-level merging paradigm. The detailed configuration of Fig. 3 is shown in Appendix B. Additional results demonstrating our findings can be found in Appendix D.
@@ 3.1.1 Stage 1: Decomposition & Initialization
Instead of treating the task vector $\tau_{\text{lrm}}$ as a monolithic update, we decompose it into orthogonal subspaces via SVD to conduct merging in the subspace level. Consider the $l$-th layer, a linear layer with a dimension $d_{\text{out}}\times d_{\text{in}}$. We apply SVD to its corresponding reasoning task vector:
@@
where $\mathbf{U}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}$ and $\mathbf{V}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ represent the left and right singular vectors, and $\mathbf{S}=\text{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ represents the singular values, indicating the merging coefficient of each subspace. We initialize the merged layers with these decomposed matrices. Crucially, to preserve the semantic direction of the reasoning task, we freeze the orthogonal bases $\mathbf{U}$ and $\mathbf{V}$ and the base intensity $\mathbf{S}$ during the entire training phase.
@@ 3.1.2 Stage 2: Injection & Training
Visual Preservation. Specifically, we treat the original base VLM $\theta_{\text{vlm}}$ as the frozen teacher and the merged model $\theta_{\text{vlr}}$ (parameterized by the learnable gates $g$) as the student. To ensure that the injection of reasoning subspaces does not disrupt the post-training vision alignment, we minimize the Kullback-Leibler (KL) divergence between the output probability distributions of the two models on a calibration dataset $\mathcal{D}$. In our experiments, we apply VizWiz (Gurari et al., 2018), a real-world visual question answering (VQA) dataset, for calibration. This objective enforces consistency in the model’s predictive behavior for visual inputs, effectively penalizing gating configurations that lead to significant semantic drift. The distillation loss is formulated as:
@@
where $P(\cdot|x;\theta)$ denotes the output probability distribution given input $x$. This objective ensures that the merged model retains the robust visual understanding capabilities of the VLM before merging.
@@
This term encourages enlarging the merging coefficients of reasoning subspaces defined by the LRM, thus improving the overall reasoning capabilities of the model. Practically, because the scales of different models vary greatly, the injection loss $\mathcal{L}_{\text{inject}}$ is normalized before training; see Appendix I. Please check Appendix G.2 for the detailed analysis of the loss terms.
@@
where $\alpha$ is a hyperparameter that controls the injection strength. This optimization process automatically discovers the subspaces that yield the optimal trade-off between reasoning injection and visual preservation. It effectively acts as a filter of subspaces: it retains reasoning components that are orthogonal to visual perception capabilities while suppressing those that cause high visual degradation (high KL divergence). It should be noted that during the training process, only the learnable gating parameters are trainable, while the original VLM weights and the decomposed singular vectors remain frozen. This ensures that the method remains highly parameter-efficient and preserves the foundational visual capabilities of the base model. The algorithm flow of the proposed FRISM is shown in the Appendix A. More technique details of FRISM is in the Appendix I.

### v2

@@ Abstract
Efficiently enhancing the reasoning capabilities of Vision-Language Models (VLMs) by merging them with Large Reasoning Models (LRMs) has emerged as a promising direction. However, existing methods typically operate at a coarse-grained layer level, which often leads to a trade-off between injecting reasoning capabilities and preserving visual capabilities. To address this limitation, we propose FRISM (Fine-grained Reasoning Injection via Subspace-level model Merging), a fine-grained reasoning injection framework based on subspace-level model merging. Observing that different SVD subspaces contribute differently to reasoning and perception, FRISM decomposes LRM task vectors via Singular Value Decomposition (SVD) and adaptively tunes the scaling coefficients of each subspace through learning to realize fine-grained reasoning injection. Furthermore, we introduce a label-free self-distillation learning strategy with dual-objective optimization using common vision-language perception datasets. Extensive experiments demonstrate that FRISM effectively improves reasoning capabilities while largely preserving the model’s visual capabilities by consistently achieving strong performance across diverse visual-language reasoning benchmarks.
Machine Learning, Model Merging, Post-Training, Multi-task Learning
@@ 1 Introduction
In recent years, represented by LLaVA (Liu et al., 2023, 2024a), Qwen-VL (Wang et al., 2024b; Bai et al., 2025), and InternVL (Chen et al., 2024b; Zhu et al., 2025), Vision Language Models (VLMs) built upon Large Language Models (LLMs) have achieved remarkable success, demonstrating strong performance across a wide range of vision language tasks, including image recognition, visual understanding, and text generation (Liu et al., 2023, 2024a; Cao et al., 2024; Brooks et al., 2023). In parallel, the emergence of large language reasoning models (LRMs) like Deepseek-R1 (DeepSeek-AI, 2025) and OpenAI-o1 (Jaech et al., 2024) has significantly advanced the frontier of model capabilities, especially in complex tasks like mathematical reasoning, problem solving, programming, and verification (Chen et al., 2025b; Zheng et al., 2025). Accordingly, many recent studies have focused on improving the reasoning capabilities of VLMs (Wang et al., 2025c), proposing various benchmarks (Yang et al., 2025b; Wang et al., 2024a) and training strategies (Wang et al., 2025a; Huang et al., 2025). These works consistently show the effectiveness of enhancing reasoning capabilities to improving the performance of VLMs.
@@
Due to the lack of labeled vision-language reasoning training datasets and the heavy post-training costs, efficiently equipping VLMs with reasoning capabilities has become an important research direction. Among existing methods (Xiao et al., 2025; Huang et al., 2025), model merging achieves this by integrating the weights of LRMs and VLMs, incurring minimal training overhead and data annotation costs. Bring Reason to Vision (BR2V) (Chen et al., 2025c) first demonstrates the feasibility of reasoning injection from LRMs to VLMs through model merging. However, common model merging techniques are developed for multi-task settings, which may compromise the visual capabilities of the merged models and cause them to fail. To this end, FRANK (Wei and Chen, 2025) proposes a layer-wise model merging method by identifying shallow and deeper layers and employing a Taylor-derived closed-form solution to guide the merging process. IP-Merging (Hu et al., 2025) only merges a subset of typical layers based on the similarity between the weights of VLMs and LRMs while leaving the remaining layers unchanged. Although effective to some extent, these methods still exhibit a significant performance gap compared to post-training-based reasoning VLMs.
@@
$W_{\text{M}}=W_{\text{base}}+\sum_{m\in\{\text{vlm},\text{lrm}\}}\lambda_{m}\cdot\tau_{m},$ (1)
where $W_{\text{M}}$ represents the weights of the merged layer, and $W_{\text{base}}$ denotes the weights of the pre-trained base model. The term $\tau_{m}$ refers to the task vector derived from model $m$, and $\lambda_{m}$ is the associated merging coefficient, where $m\in\{\text{vlm},\text{lrm}\}$ corresponds to VLM and LRM, respectively. Fig. 1 further shows the diagram of (a) individual models, (b) layer-wise merging. For individual models, $\lambda_{\text{vlm}}=1,\lambda_{\text{lrm}}=0$ recovers the original VLM, while $\lambda_{\text{vlm}}=0,\lambda_{\text{lrm}}=1$ recovers the original LRM. Layer-wise merging methods assign layer-dependent values to $\lambda_{\text{vlm}}$ and $\lambda_{\text{lrm}}$, where a larger $\lambda_{\text{vlm}}$ indicates stronger visual capability and a larger $\lambda_{\text{lrm}}$ indicates stronger reasoning capability. However, such layer-wise balancing is inherently coarse-grained and can only produce a compromise: tuning $\lambda_{\text{vlm}}$ and $\lambda_{\text{lrm}}$ at the layer level entangles the two capabilities, leading to unavoidable losses in either vision or reasoning capabilities. Fig. 2 shows the average performance on two vision-language (VL) reasoning datasets, MMStar (Chen et al., 2024a) and R1-OneVision (Yang et al., 2025b), and two VL perception datasets, TextVQA (Singh et al., 2019) and POPE (Li et al., 2023b). The results show that existing methods usually lead to a trade-off between VL reasoning and perception performance.
In this paper, we observe that a layer is not the atomic unit of model capability: even within a single layer, reasoning capabilities and visual representations can reside in distinct subspaces. Motivated by the low-rank nature and the compact encoding of high-level or reasoning-related functions of parameter updates in LLMs (Cai et al., 2025; Ping et al., 2024; Zhou et al., 2025), we use the subspaces of the LRM task vectors obtained through singular value decomposition (SVD) as the reasoning prior and tune the injection magnitude of each subspace. We find that the gains from different ranks peak at different magnitudes. This heterogeneity suggests that a uniform layer-wise scaling coefficient is insufficient: it inevitably couples beneficial reasoning, which may require high magnitude, with visual noise, thereby limiting overall improvements.
Thus, to achieve the maximum reasoning capabilities with the minimum degree of vision loss, we propose FRISM (Fine-grained Reasoning Injection via Subspace-level model Merging). As shown in Fig. 1 (c), we change the merging paradigm to:
@@
  • We rethink the paradigm of merging VLMs and LRMs to efficiently enhance visual reasoning capabilities. Instead of coarse-grained merging at the layer-level, we propose a fine-grained merging paradigm to realize more effective reasoning injection.
  • We first propose a subspace-level model merging for reasoning injection from LRMs to VLMs. By decomposing LRM task vectors via SVD and modulating at the subspace granularity, our method enables the adaptive reweighting of SVD subspaces under a visual-preservation objective.
  • We design label-free self-distillation training to relieve the scarcity of multimodal reasoning data. Through a dual-objective optimization that safeguards visual consistency while maximizing spectral injection intensity, our approach achieves stable and efficient reasoning transfer without requiring VL reasoning datasets.
@@ 2 Related Work
Model Merging integrates models finetuned on different tasks within the parameter space to assimilate diverse capabilities (Yang et al., 2024b). In multi-task settings, the critical challenge for model merging is weight interference. Various methods are proposed to address this issue, including calculating information matrices (Jin et al., 2022; Matena and Raffel, 2022), merging coefficient tuning (Ilharco et al., 2023; Yang et al., 2023), task vector pruning (Yadav et al., 2023; Yu et al., 2024; Huang et al., 2024; Gargiulo et al., 2025; Marczak et al., 2025), and dynamic routing (Ye et al., 2025; Tang et al., 2024; Lu et al., 2024b; Shen et al., 2024; Yang et al., 2024c), Recently, this paradigm has been extended to merge VLMs and LRMs (Chen et al., 2025c), aiming to transfer the reasoning capabilities between models trained on different modalities. However, methods designed under multi-task settings often incur severe performance loss. To relieve this issue, FRANK (Wei and Chen, 2025) performs layer-wise weighted fusion based on Taylor expansion and IP-Merging (Hu et al., 2025) mitigates cross-modal conflicts via parameter projection based on weight similarity. Nevertheless, limited by coarse granularity or static thresholding, these methods struggle to distinguish parameters beneficial for reasoning from those that cause visual interference. Therefore, effective Reasoning-to-Vision transfer requires not merely conflict mitigation but also precise control over task vectors. To this end, we propose FRISM, which fine-grainedly localizes and preserves parameter components that are critical for reasoning without harming visual capabilities.
Subspace Decomposition technique aims to factorize high-dimensional matrices to reveal latent structures. Methods such as Non-negative Matrix Factorization (NMF) and Independent Component Analysis (ICA) demonstrate advantages in interpretability or disentanglement. Among them, SVD serves as a fundamental subspace decomposition technique offering optimal low-rank approximations. It stands as a cornerstone in deep learning for model compression (Idelbayev and Carreira-Perpinán, 2020) and parameter-efficient fine-tuning (PEFT) (Hu et al., 2022; Meng et al., 2024). Recent studies (Cai et al., 2025; Sharma et al., 2024; Zhang et al., 2023; Ping et al., 2024; Yang et al., 2025a) indicate that model capabilities are not uniformly distributed but are distinctively encoded within specific rank subspaces. Motivated by these findings, we treat LRM task vectors as inference priors. By recalibrating their spectral distribution via SVD, FRISM seamlessly adapts textual reasoning priors to the VL domain, thereby enhancing VLM reasoning capabilities without compromising visual performance.
@@ 3 Method
where $\tau$ are task vectors. Existing methods usually conduct model-level merging (Chen et al., 2025c):
$\theta_{\text{vlrm}}=\theta_{\text{base}}+\lambda_{\text{vlm}}\cdot\tau_{\text{vlm}}+\lambda_{\text{lrm}}\cdot\tau_{\text{lrm}},$ (4)
@@
$\theta_{\text{vlrm}}^{(l)}=\theta_{\text{base}}^{(l)}+\lambda_{\text{vlm}}^{(l)}\cdot\tau_{\text{vlm}}^{(l)}+\lambda_{\text{lrm}}^{(l)}\cdot\tau_{\text{lrm}}^{(l)},\quad l=1..L$ (5)
@@
Motivation. We hypothesize that reasoning capabilities and visual representations reside in distinct subspaces within the parameter space. Consequently, a coarse-grained, layer-wise merging coefficient may be insufficient to disentangle these capabilities. To validate this hypothesis, we analyze the impact of merging different subspaces of the LRM task vector. Specifically, we decompose the task vector of DeepSeek-R1-Distill-Qwen-7B, via SVD and inject specific rank subspaces into the VLM, Qwen2.5-VL-Instruct, with varying scaling coefficients. We observe that (1) different subspaces achieve peak performance at distinctly different scaling coefficients; (2) standard layer-wise merging results in suboptimal performance. These observations suggest that subspace-level modulation is a useful granularity for reasoning injection. Therefore, to achieve maximum reasoning capabilities with minimal vision loss, it is essential to shift from coarse-grained layer-level merging to a fine-grained subspace-level merging paradigm. The detailed configuration of Fig. 3 is shown in Appendix B. Additional results demonstrating our findings can be found in Appendix D.3.
@@ 3.1.1 Stage 1: Decomposition & Initialization
Instead of treating the task vector $\tau_{\text{lrm}}$ as a monolithic update, we decompose it into orthogonal subspaces via SVD to conduct merging at the subspace level. Consider the $l$-th layer, a linear layer with a dimension $d_{\text{out}}\times d_{\text{in}}$. We apply SVD to its corresponding reasoning task vector:
@@
where $\mathbf{U}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{V}\in\mathbb{R}^{r\times d_{\text{in}}}$ represent the left and right singular vectors, and $\mathbf{S}=\text{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{r\times r}$ represents the singular values, indicating the merging coefficient of each subspace. We initialize the merged layers with these decomposed matrices. Crucially, to preserve the semantic direction of the reasoning task, we freeze the orthogonal bases $\mathbf{U}$ and $\mathbf{V}$ and the base intensity $\mathbf{S}$ during the entire training phase.
@@ 3.1.2 Stage 2: Injection & Training
Visual Preservation. Specifically, we treat the original base VLM $\theta_{\text{vlm}}$ as the frozen teacher and the merged model $\theta_{\text{vlrm}}$ (parameterized by the learnable gates $g$) as the student. To ensure that the injection of reasoning subspaces does not disrupt the post-training vision alignment, we minimize the Kullback-Leibler (KL) divergence between the output probability distributions of the two models on a calibration dataset $\mathcal{D}$. In our experiments, we apply VizWiz (Gurari et al., 2018), a real-world visual question answering (VQA) dataset, for calibration. This objective enforces consistency in the model’s predictive behavior for visual inputs, effectively penalizing gating configurations that lead to significant semantic drift. The distillation loss is formulated as:
@@
where $P(\cdot|x;\theta)$ denotes the output probability distribution given input $x$. This objective encourages the merged model to retain the robust visual understanding capabilities of the VLM before merging.
@@
This term encourages enlarging the merging coefficients of reasoning subspaces defined by the LRM, thus improving the overall reasoning capabilities of the model. Practically, because the scales of different models vary greatly, the injection loss $\mathcal{L}_{\text{inject}}$ is normalized before training; see Appendix H. Please check Appendix F.2 for the detailed analysis of the loss terms.
@@
where $\alpha$ is a hyperparameter that controls the injection strength. This optimization process automatically discovers the subspaces that yield the optimal trade-off between reasoning injection and visual preservation. It effectively acts as a filter of subspaces: it retains reasoning components that are orthogonal to visual perception capabilities while suppressing those that cause high visual degradation (high KL divergence). It should be noted that during the training process, only the learnable gating parameters are trainable, while the original VLM weights and the decomposed singular vectors remain frozen. This ensures that the method remains highly parameter-efficient and preserves the foundational visual capabilities of the base model. The algorithm flow of the proposed FRISM is shown in the Appendix A. More technique details of FRISM is in the Appendix H.